algebraically, find the intersection points of the two lines given in each part of this problem. y=0.12x+18 y=0.08x+25
1
Expert's answer
2012-08-10T10:44:01-0400
To find intersection point we need to equate right parts of given expressions:
0.12x+18=0.08x+25
0.12x-0.08x=25-18
0.04x=7
x=7/0.04=175
To find corresponding value of y we should substitute obtained x into any of the expressions:
y=0.12*175+18=21+18=39
So the intersection point is (175,39)